OpenStudy (anonymous):

Solve for t. t^6+1t^3=20 It's not right but I got -5,4 and I can't figure out what I did wrong.

OpenStudy (anonymous):

t is 4 i think

OpenStudy (anonymous):

but i suck at math so i might be wrong

jimthompson5910 (jim_thompson5910):

t^6+t^3=20 t^6+t^3-20 = 0 (t^3)^2+t^3-20 = 0 Let z = t^3. So z^2 = (t^3)^2 z^2 + z - 20 = 0

jimthompson5910 (jim_thompson5910):

So if you solve z^2 + z - 20 = 0 for z, you get z = 4 and z = -5

jimthompson5910 (jim_thompson5910):

Your last step is to solve t^3 = 4 and t^3 = -5

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

yw